# Computational Modeling for Personalized Transcranial Electrical Stimulation: Theory, Tools, and Applications

### v1

@@ Abstract
Objective. Personalized transcranial electrical stimulation (tES) has gained growing attention due to the substantial inter-individual variability in brain anatomy and physiology. While previous reviews have discussed the physiological mechanisms and clinical applications of tES, there remains a critical gap in up-to-date syntheses focused on the computational modeling frameworks that enable individualized stimulation optimization. Approach. This review presents a comprehensive overview of recent advances in computational techniques supporting personalized tES. We systematically examine developments in forward modeling for simulating individualized electric fields, as well as inverse modeling approaches for optimizing stimulation parameters. We critically evaluate progress in head modeling pipelines, optimization algorithms, and the integration of multimodal brain data. Main results. Recent advances have substantially accelerated the construction of subject-specific head conductor models and expanded the landscape of optimization methods, including multi-objective optimization and brain network-informed optimization. These advances allow for dynamic and individualized stimulation planning, moving beyond empirical trial-and-error approaches. Significance. By integrating the latest developments in computational modeling for personalized tES, this review highlights current challenges, emerging opportunities, and future directions for achieving precision neuromodulation in both research and clinical contexts.
@@ 1 Introduction
Transcranial electrical stimulation (tES) is a non-invasive neuromodulation technique that delivers low-intensity electrical currents to the scalp with the objective of modulating brain activity. Compared to invasive techniques such as deep brain stimulation (DBS), tES offers great advantages in safety, accessibility, and cost-effectiveness [1, 2]. Accumulating evidence indicates that tES can modulate neural oscillations and cortical excitability, thereby influencing a range of cognitive functions and behaviors [3, 4, 5, 6]. As such, tES has been widely applied as both a research tool to probe causal brain-behavior relationships and a therapeutic intervention for neurological and psychiatric conditions [7, 8, 9, 10, 11, 12].
The efficacy of tES is influenced by various factors, including electric field strength and orientation, stimulation frequency, waveform characteristics, and even the ongoing brain state [1, 3, 13]. Multiple tES modalities have been developed, including transcranial direct current stimulation (tDCS) [14, 15], transcranial alternating current stimulation (tACS) [16, 17], transcranial random noise stimulation (tRNS) [18], intersectional short pulse stimulation (ISP) [19], and temporal interference stimulation (TIS) [20, 21, 22]. Each modality offers distinct mechanisms of action and application potentials, providing researchers and clinicians with versatile tools for neuromodulation.
A growing body of research has highlighted the substantial inter-individual variability in neuroanatomy and neurophysiology, prompting increasing interest in personalized tES strategies [23, 24, 25, 22]. While several earlier works touched upon personalization in tES [26, 27, 28], most reviews have emphasized physiological mechanisms or clinical efficacy [3, 29, 30], with comparatively less attention paid to the computational frameworks that support individualized targeting and optimization. Given that computational models now play a central role not only in estimating electric field distributions but also in systematically optimizing stimulation parameters, a focused synthesis of these advances is urgently needed.
Computational models in tES serve a dual role: forward simulation (Fig.1 A-E) and inverse optimization (Fig.1 F-I). Forward modeling, often referred to as head volume conduction modeling, predicts the spatial distribution of electric fields induced within the brain during stimulation. This capability is essential for mechanistically understanding how tES modulates neural activity [31]. Traditional forward models often rely on standardized anatomical templates such as MNI152 [32, 33] or ICBM152 [34]. However, substantial inter-individual variability in head anatomy, tissue conductivity, and cerebrospinal fluid distribution can lead to marked differences in electric field profiles even under identical stimulation conditions [35, 36, 37]. Consequently, constructing subject-specific head models from individual MRI scans has become critical for achieving accurate simulations. Recent advances in neuroimaging acquisition and segmentation pipelines have dramatically reduced the time required to build individualized head models, paving the way for broader clinical and research applications.
Alongside the evolution of modeling techniques, tES hardware has also advanced—from simple two-electrode setups to complex multi-electrode arrays—substantially expanding the parameter space for stimulation configuration. Manual montage design or simple forward simulation is often insufficient to identify optimal stimulation strategies in this high-dimensional space [38]. Inverse optimization methods have thus emerged as essential tools, integrating personalized forward models with optimization algorithms to systematically determine stimulation parameters that maximize efficacy, enhance target specificity, and minimize unintended effects. Optimization techniques explored in the field include exhaustive search [39, 40], quadratic programming [41], gradient-based methods [42], and heuristic algorithms such as genetic optimization [22, 43]. Critical factors in inverse modeling include the definition of biologically meaningful objective functions, adherence to safety constraints, and accommodation of hardware limitations. Moreover, integrating multimodal information, such as functional or structural connectivity data, offers opportunities for network-informed dynamic stimulation designs. These recent advances and their implications for future personalized tES development are systematically discussed in this review.
This review aims to fill this gap by providing a comprehensive and structured overview of computational modeling approaches for personalized tES. Specifically, we examine forward modeling (used to predict electric field distributions) in Section 2 and inverse optimization (used to tailor stimulation strategies) in Section 3, and discuss how recent advances in neuroimaging, numerical solvers, optimization algorithms, and multimodal data integration are shaping the future of precision neuromodulation in Section 4. Finally, we discuss the current limitations of computational modeling in tES and outline promising future directions for research and application.
@@
Figure 1: Overview of the forward modeling and inverse optimization for personalized tES. (A) The construction of forward models for personalized computational simulations begins with tissue segmentation from MRI data. (B) Conductivity values are assigned to each segmented tissue type to generate an individualized computational model. (C) Head conductor model is generated by mesh generation and numerical methods. (D) Virtual electrodes are then positioned on the model, and current is applied according to the predefined stimulation protocol. (E) The electric field distribution is simulated by solving the current flow within the head model. The stimulation montage is optimized based on the electric field simulation results from the forward model. (F) The stimulation target is first identified, followed by (G) verifying current safety constraints. (H) An optimization algorithm is then applied to the model to determine optimal stimulation parameters. (I) Based on the optimization results, a stimulation montage is applied to the personalized tES for the subject.
@@ 2 Forward modeling for personalized tES
Forward modeling, also known as volume conduction modeling, plays a fundamental role in simulating the distribution of electric fields generated by (Table LABEL:tab:headmodel) (Fig.1 A-E). The earliest modeling work employed simplified spherical head models, which approximated the human head as three nested concentric spheres representing the scalp, skull, and brain, respectively. Despite their simplicity, these models effectively captured key aspects of electric field distribution and demonstrated reasonable agreement with experimental measurements [44]. As the field progressed and the demand for realistic head models increased, researchers began incorporating real anatomical data into head conductivity models using MRI-based imaging. The construction of realistic forward models typically involves several major steps: tissue segmentation, mesh generation, and electric field simulation. In the early stages, constructing individualized models was time-consuming and labor-intensive, prompting the widespread use of standardized templates for forward modeling. Among the most prominent templates are the MNI152 head model [32, 33], derived from the average of 152 anatomical scans, and the New York Head Model introduced by Huang et al.[45], a finite element method (FEM)-based model built upon the ICBM152 template. These resources have significantly advanced large-scale and reproducible tES research [46, 47, 48]. Recent advances in neuroimaging pipelines and high-performance computing have made the construction of subject-specific head models increasingly practical. Personalized head models can now be generated from individual MRI within a few hours, often fully automatically, thus enabling individualized simulation of electric fields prior to tES intervention.
In the following sections, we review the formulation and evolution of forward modeling approaches for personalized tES, from early spherical models to anatomically realistic, individualized simulations. This progression illustrates the field’s continuous efforts to improve the accuracy and targeting precision of tES through increasingly advanced computational tools.
@@ 2.1 Tissue Segmentation, Electrode Co-registration, Mesh Generation and Numerical Methods
Segmentation involves delineating key tissues relevant to electrical conductivity, including the scalp, skull, cerebrospinal fluid (CSF), gray matter (GM), and white matter (WM). Conductivity values derived from empirical measurements are assigned to segmented tissues to enable electric field distribution calculations (Fig. 1 B).Accurate segmentation is the first critical step toward effective modeling. Segmentation quality is commonly assessed using metrics such as the Dice coefficient and the modified Hausdorff distance, each providing complementary insights. The Dice coefficient measures the similarity between two sets, by computing the ratio of twice their intersection to the sum of their cardinalities. While the Dice coefficient effectively captures the overlap of large voxel clusters, it may overlook subtle discrepancies along region boundaries. To address this limitation, the modified Hausdorff distance [49] is often employed. It calculates the average of two directed Hausdorff distances, where each directed distance represents the maximum deviation from a point in one set to its nearest neighbor in the other. By focusing on boundary dissimilarities, the modified Hausdorff distance offers a finer-grained evaluation of segmentation accuracy. In practice, better segmentation results are indicated by a higher Dice coefficient and a lower modified Hausdorff distance.
@@
The Finite Element Method (FEM) is the most widely adopted [50, 51], owing to its flexibility in handling complex geometries and heterogeneous, anisotropic conductivity distributions [52]. FEM discretizes the volume into small elements, enabling high-fidelity approximations of electric fields in intricate anatomical structures. An alternative approach is the Boundary Element Method (BEM), which simplifies the problem by meshing only the boundaries between tissues [53, 54]. By reducing the problem dimensionality, BEM substantially lowers computational demands. However, its applicability is limited in cases involving significant conductivity inhomogeneity or anisotropy [55, 56]. Despite these challenges, recent methodological advances have improved BEM’s robustness for tES applications [57, 58]. The Finite Difference Method (FDM) presents another option, discretizing the domain on structured grids. While FDM is conceptually simple and computationally efficient for regular geometries, it struggles with irregular anatomical boundaries and spatially varying conductivities, which can compromise simulation accuracy [59].
2.2 Finite Element Method for Electrical Field Simulation
The FEM, due to its flexibility, is well-suited for handling complex geometries and is most commonly used for modeling and solving the electric field distribution within the brain. Here, we will use the FEM as an example for a detailed introduction. Once the whole head mesh model in generated The accuracy of these simulations critically depends on the quality of the preceding tissue segmentation and mesh generation steps (Fig. 1 D). The meshed head model incorporates the spatial distribution of conductivity across different tissues, and the electric field simulation is governed by Laplace’s equation (Fig. 1 E). Assuming the absence of net current sources or sinks within the volume, the divergence of the current density ($J$) vanishes. When external currents are applied through electrodes placed on the scalp, the resulting potential distribution ($V$) within the head can be computed by solving the following form of Laplace’s equation [60]:
@@ 2.2 Finite Element Method for Electrical Field Simulation
After assembling the global matrix and applying the boundary conditions, the system is solved numerically to obtain the distribution of the electric field. To efficiently compute the potential distribution, open-source solvers such as getDP [61] can be utilized.
Table 1: Head model (volume conduction model)
Reference Meshing Methods Models
Miranda P C, et al. (2006). Clinical neurophysiology [44] FEM Spherical Head Model
Wagner T A, et al. (2004). IEEE Transactions on Biomedical Engineering [51] FEM Realistic Human Head Model
Datta A, et al. (2009). Brain stimulation [62] FEM Realistic Human Head Model
Rampersad S M, et al. (2014). IEEE Transactions on Neural Systems and Rehabilitation Engineering [63] FEM Realistic Human Head Model
Fuchs M, et al. (2002). Clinical neurophysiology [54] BEM Realistic Human Head Model
Susnjara A, et al. (2019). International Conference on Software, Telecommunications and Computer Networks (SoftCOM) [57] BEM Realistic Human Head Model
Šušnjara A, et al. (2022). Engineering Analysis with Boundary Elements [58] BEM Realistic Human Head Model
Magsood, H., et al. (2021). Materials Science and Engineering [64] FDM Realistic Human Head Model
Gomez, L. J., et al. (2022). Brain stimulation [65] FEM, BEM, FDM Realistic Human Head Model
2.3 Simulating electrical field distribution induced by tES
After obtaining the electric field in a single direction by anode and cathode, according to [66], the averaged electric field norm of tDCS can be formulated as Eq. (3).

### v2

# Personalized Transcranial Electrical Stimulation: A Review of Computational Modeling and Optimization

@@ Abstract
Objective. Personalized transcranial electrical stimulation (tES) has gained increasing attention due to the substantial inter-individual variability in brain anatomy and physiology. While previous reviews have discussed the physiological mechanisms and clinical applications of tES, there remains a critical gap in up-to-date syntheses focused on the computational modeling frameworks that enable individualized stimulation optimization. Approach. This review presents a comprehensive overview of recent advances in computational techniques supporting personalized tES. We systematically examine developments in forward modeling for simulating individualized electric fields, as well as inverse modeling approaches for optimizing stimulation parameters. We critically evaluate progress in head modeling pipelines, optimization algorithms, and the integration of multimodal brain data. Main results. Recent advances have substantially accelerated the construction of subject-specific head conductor models and expanded the landscape of optimization methods, including multi-objective optimization and brain network-informed optimization. These advances allow for dynamic and individualized stimulation planning, moving beyond empirical trial-and-error approaches. Significance. By integrating the latest developments in computational modeling for personalized tES, this review highlights current challenges, emerging opportunities, and future directions for achieving precision neuromodulation in both research and clinical contexts.
@@ 1 Introduction
Transcranial electrical stimulation (tES) is a non-invasive neuromodulation technique that delivers low-intensity electrical currents to the scalp with the objective of modulating brain activity. Compared to invasive techniques such as deep brain stimulation (DBS), tES offers great advantages in safety, accessibility, and cost-effectiveness [Krause2023TranscranialES, huang2017measurements]. Accumulating evidence indicates that tES can modulate neural oscillations and cortical excitability, thereby influencing a range of cognitive functions and behaviors [liu2018immediate, Maiella2022SimultaneousTE, rufener2016transcranial, moliadze2019after]. As such, tES has been widely applied as both a research tool to probe causal brain-behavior relationships and a therapeutic intervention for neurological and psychiatric conditions [tavakoli2017transcranial, seo2019relation, solanki2021investigating, rasmussen2021high, simula2022transcranial, zhang2022alpha].
The efficacy of tES is influenced by various factors, including electric field strength [gaugain2023quasi, laakso2019can] and orientation [shahid2015use], stimulation frequency, waveform characteristics [louviot2022transcranial], and even the ongoing brain state [Krause2023TranscranialES, liu2018immediate, krause2019transcranial]. Multiple tES variations have been developed, including transcranial direct current stimulation (tDCS) [Nitsche2000ExcitabilityCI, Brunoni2012ClinicalRW], transcranial alternating current stimulation (tACS) [Antal2008ComparativelyWA, Bchinger2017ConcurrentTR], transcranial random noise stimulation (tRNS) [vanderGroen2016TranscranialRN], intersectional short pulse stimulation (ISP) [voroslakos2018direct], and transcranial temporal interference stimulation (tTIS) [grossman2017noninvasive, violante2022non, wang2022multi]. Because each variation has distinct mechanisms of action and application potential, computational models are essential tools for their investigation. Computational models in tES serve a dual role: forward simulation (Fig.1 A-E) and inverse optimization (Fig.1 F-I). Forward modeling, often referred to as head volume conduction modeling, predicts the spatial distribution of electric fields induced within the brain during stimulation. This capability is essential for mechanistically understanding how tES modulates neural activity [lee2021future]. In early studies, researchers used template head models (e.g., MNI152 template) to simulate the electric field. While computationally efficient, such templates can overlook individual anatomical and physiological variability, reducing predictive accuracy. Individualized models, while demanding high-quality imaging and advanced segmentation, enable precise electric field estimation and tailored stimulation optimization. Consequently, constructing subject-specific head models from individual MRI scans has become critical for achieving accurate simulations. Alongside the evolution of forward modeling techniques, tES hardware has also advanced—from simple two-electrode setups to complex multi-electrode arrays—substantially expanding the parameter space for stimulation configuration. Manual montage design or simple forward simulation is often insufficient to identify optimal stimulation strategies in this high-dimensional parameter space [hunold2023review]. Inverse modeling has thus emerged as an essential toolset: it has been widely applied in EEG and ECoG source localization to reconstruct neural current sources from measured scalp potentials, and it has been adapted to tES for optimizing electrode configurations [grech2008review, baillet2002electromagnetic, michel2004eeg]. By integrating personalized forward models with optimization algorithms, inverse modeling systematically determines individualized tES strategies that maximize efficacy, enhance target specificity, and minimize unintended effects.
A growing body of work demonstrates substantial inter-individual variability in neuroanatomy and neurophysiology, motivating increasing interest in personalized tES strategies [wang2022multi, violante2023non, ma2024mapping, grover2022long]. Although several prior articles discuss personalization in tES, most reviews primarily emphasize physiological mechanisms or clinical efficacy [liu2018immediate, yavari2018basic, chang2022application, gomez2024perspectives]. By contrast, surveys on computational modeling often provide limited, practice-oriented guidance for constructing forward models and public pipelines [gomez2024perspectives, berger2025human, bikson2012computational, van2023outcome, fernandez2020unification]; several omit optimization algorithms [hunold2023review, berger2025human, bikson2012computational, van2023outcome, bai2013review] and others do not cover tTIS or personalized modeling [salvador2025group, miranda2018realistic].
In this review, we place greater emphasis on the evolution of computational modeling itself, providing a systematic perspective on how personalized tES simulation and optimization can be implemented from the ground up. We further discuss potential enhancement techniques, including forward-model refinement through segmentation, conductivity, and anisotropy updates, neural-dynamics-integrated personalized modeling, and brain-state-informed inverse optimization. Specifically, Section 2 focuses on forward modeling for predicting electric field distributions, while Section 3 examines inverse optimization for designing personalized stimulation strategies. These sections include 64 key studies: 22 modeling and theoretical works (e.g., head model construction, electric field computation, conductivity modeling), 34 studies on stimulation optimization (covering traditional algorithms, machine learning–based methods, and individualized electrode layout design), and 8 studies involving model validation or experimental verification. Section 4 introduces advanced techniques to enhance tES modeling, incorporating 17 studies with 12 proposed frameworks and 5 experimental validation. Section 5 examines current limitations and future perspectives, particularly in relation to the integration of neuroimaging techniques, and synthesizes insights from 13 recent innovative studies. Finally, we highlight the current limitations of computational modeling in tES and outline potential directions for future research and applications.
@@
Figure 1: Overview of the forward modeling and inverse optimization for personalized tES. (A) The construction of forward models for personalized computational simulations begins with tissue segmentation from MRI data. (B) Conductivity values are assigned to each segmented tissue type to generate an individualized computational model. (C) Head conductor model is generated by mesh generation and numerical methods. (D) Virtual electrodes are then positioned on the model, and current is applied according to the predefined stimulation protocol. (E) The electric field distribution is simulated by solving the current flow within the head model. The stimulation montage is optimized based on the electric field simulation results from the forward model. (F) The stimulation target is first identified, followed by (G) verifying current safety constraints. (H) An optimization algorithm is then applied to the model to determine optimal stimulation parameters. (I) Based on the optimization results, a stimulation montage is applied to the personalized tES for the subject. Steps (F–I) illustrate how the stimulation target, safety constraints, and optimization algorithms can use forward model results to determine optimal stimulation parameters. In this section, the focus remains on forward modeling, while inverse optimization is briefly illustrated to provide context on the overall computational workflow.
@@ 2 Forward modeling for personalized tES
Forward modeling, also known as volume conduction modeling, plays a fundamental role in simulating the distribution of electric fields generated by tES (Fig.1 A-E). The earliest modeling work employed simplified spherical head models, which approximated the human head as three nested concentric spheres representing the scalp, skull, and brain, respectively. Despite their simplicity, these models effectively captured key aspects of electric field distribution, such as the spatial variation of field intensity across tissue layers and the general orientation of induced fields [rush2008eeg], and demonstrated reasonable agreement with experimental measurements, including scalp electrode recordings and intracranial measurements in in vivo studies [miranda2006modeling]. As the field progressed and the demand for realistic head models increased, researchers began incorporating real anatomical data into head conductivity models using MRI-based imaging. The construction of realistic forward models typically involves several major steps: tissue segmentation, mesh generation, and electric field simulation. In the early stages, constructing individualized models was time-consuming and labor-intensive, prompting the widespread use of standardized templates for forward modeling. Among the most prominent templates is the MNI152 template head [evans19933d, fonov2011unbiased], which represents the average of 152 T1-weighted anatomical scans. FEM-based TES models, such as the New York Head Model introduced by Huang et al. [huang2016new], are built upon such anatomical templates and incorporate additional information, including tissue conductivity, to simulate electric fields. These resources have significantly advanced large-scale and reproducible tES research [asadzadeh2020systematic, salehinejad2020transcranial, zorzos2021advances]. However, for normalized models, substantial inter-individual variability in head anatomy and tissue conductivity, which can lead to marked differences in electric field profiles even under identical stimulation conditions [kasten2019integrating, von2021interindividual, antonenko2021inter]. By contrast, individualized models are based on subject-specific MRI data, allowing accurate reconstruction of anatomical details and tissue properties. While they require high-quality imaging and advanced segmentation, these models enable subject-tailored optimization of stimulation parameters, thereby providing more precise electric field predictions and greater suitability for clinical applications. Recent advances in neuroimaging pipelines and high-performance computing have made the construction of subject-specific head models increasingly practical. Personalized head models can now be generated from individual MRI within a few hours, often fully automatically, thus enabling individualized simulation of electric fields prior to tES intervention.
In the following sections, we review the formulation and evolution of forward modeling approaches for personalized tES. Section 2.1 covers tissue segmentation, electrode alignment, and mesh generation. Section 2.2 introduces the finite element method for simulating electric fields. Section 2.3 summarizes approaches for simulating electric field distributions induced by tES. Section 2.4 highlights software platforms that enable individualized head model construction. This section provides a theoretical and technical foundation for precise modeling and application of personalized tES.
@@ 2.1 Tissue Segmentation, Electrode Co-registration, Mesh Generation and Numerical Methods
Segmentation involves delineating key tissues relevant to electrical conductivity, including the scalp, skull, cerebrospinal fluid (CSF), gray matter (GM), and white matter (WM). Conductivity values derived from empirical measurements are assigned to segmented tissues to enable electric field distribution calculations (Fig. 1 B).Accurate segmentation is the first critical step toward effective modeling. Segmentation quality is commonly assessed using metrics such as the Dice coefficient and the modified Hausdorff distance, each providing complementary insights. The Dice coefficient measures the similarity between two sets, by computing the ratio of twice their intersection to the sum of their cardinalities. While the Dice coefficient effectively captures the overlap of large voxel clusters, it may overlook subtle discrepancies along region boundaries. To address this limitation, the modified Hausdorff distance [dubuisson1994modified] is often employed. It calculates the average of two directed Hausdorff distances, where each directed distance represents the maximum deviation from a point in one set to its nearest neighbor in the other. By focusing on boundary dissimilarities, the modified Hausdorff distance offers a finer-grained evaluation of segmentation accuracy. In practice, better segmentation results are indicated by a higher Dice coefficient and a lower modified Hausdorff distance.
@@
The Finite Element Method (FEM) is the most widely adopted [miranda2006modeling, datta2009gyri, khalil1977parametric, wagner2004three, rampersad2014simulating], owing to its flexibility in handling complex geometries and heterogeneous, anisotropic conductivity distributions [wagner2007transcranial]. FEM discretizes the volume into small elements, enabling high-fidelity approximations of electric fields in intricate anatomical structures. An alternative approach is the Boundary Element Method (BEM), which simplifies the problem by meshing only the boundaries between tissues [salinas20093d, susnjara2019stochastic, vsuvsnjara2022uncertainty]. By reducing the problem dimensionality, BEM substantially lowers computational demands. However, its applicability is limited in cases involving significant conductivity inhomogeneity or anisotropy [dayarian2023evaluating, vsuvsnjara2021stochastic]. Despite these challenges, recent methodological advances have improved BEM’s robustness for tES applications [susnjara2019stochastic, vsuvsnjara2022uncertainty]. The Finite Difference Method (FDM) presents another option, discretizing the domain on structured grids. While FDM is conceptually simple and computationally efficient for regular geometries, it may face limitations when applied to irregular anatomical boundaries and spatially varying conductivities [turovets20143d]. However, modern numerical (e.g., MagStim) schemes and some commercial software have implemented strategies that mitigate these challenges, for example through refined discretization and dedicated treatments of non-rectangular geometries. Consequently, FDM-based solvers can still provide accurate solutions for many practical head models.
2.2 Finite Element Method for Electric Field Simulation
The FEM, due to its flexibility, is well-suited for handling complex geometries and is most commonly used for modeling and solving the electric field distribution within the brain. Here, we will use the FEM as an example for a detailed introduction. Once the whole head mesh model is generated, the accuracy of these simulations critically depends on the quality of the preceding tissue segmentation and mesh generation steps (Fig. 1 D). The meshed head model incorporates the spatial distribution of conductivity across different tissues, and the electric field simulation is governed by Laplace’s equation (Fig. 1 E). Assuming the absence of net current sources or sinks within the volume, the divergence of the current density ($J$) vanishes. When external currents are applied through electrodes placed on the scalp, the resulting potential distribution ($V$) within the head can be computed by solving the following form of Laplace’s equation [dmochowski2011optimized]:
@@ 2.2 Finite Element Method for Electric Field Simulation
After assembling the global matrix and applying the boundary conditions, the system is solved numerically to obtain the distribution of the electric field. To efficiently compute the potential distribution, open-source solvers such as getDP [dular1998general] can be utilized.
2.3 Simulating Electric Field distribution induced by tES
After obtaining the electric field in a single direction by anode and cathode, according to [saturnino2019accessibility], the averaged electric field norm of tDCS can be formulated as Eq. (3).